Business Weekly conducted a survey of graduates from 30 top MBA programs. On the basis of the survey, assume the mean annual salary for graduates 10 years after graduation is $157,000. Assume the standard deviation is $33,000. Suppose you take a simple random sample of 44 graduates. Round all answers to four decimal places if necessary.

What is the distribution of XX? XX ~ N(,)

What is the distribution of ¯xx¯? ¯xx¯ ~ N(,)

For a single randomly selected graduate, find the probability that her salary is between $155,438 and $161,425.

For a simple random sample of 44 graduates, find the probability that the average salary is between $155,438 and $161,425.

For part d), is the assumption of normal necessary? NoYes
Answer #1

1)

distribution of X ~N(157000,33000)

2)distribution of ¯x ~ N(157000, 4974.9372)

3)

probability =P(155438<X<161425)=P((155438-157000)/33000)<Z<(161425-157000)/33000)=P(-0.05<Z<0.13)=0.5533-0.4811=0.0722

4)

probability =P(155438<X<161425)=P((155438-157000)/4974.937)<Z<(161425-157000)/4974.937)=P(-0.31<Z<0.89)=0.8131-0.3768=0.4364

5)No since sample size >30